OpenStudy (anonymous):

confirm that f and g are inverses by showing that f(g(x))=x and g(f(x)) = x . f(x)=4/x and g(x) = 4/x .

OpenStudy (lgbasallote):

g(x) = 4/x therefore f(g(x)) = \(\large \frac{4}{\frac{4}{x}} \implies 4 \times \frac x4 \implies x\) got it?

OpenStudy (anonymous):

yep i got it !thankyou

OpenStudy (lgbasallote):

welcome welcome ^_^